OpenStudy (anonymous):

Express these numbers in terms of i. -√-9

OpenStudy (anonymous):

-3i

OpenStudy (anonymous):

-3i

OpenStudy (hoblos):

i^2 = -1 so -9 = 9i^2 -√-9 = -√(9i^2) = -3i